Meeting Information

  • 1600 UT (1000 MDT)
  • AOC 317; IP:; phone: +1 575-835-7370


News, Travel, Visitors (McMullin)

  • Overall status: In B configuration. Some antenna specific issues (EA22 pointing) which have been resolved over several days. No observing last night.
  • Key dates:
    • 09-23 Feb: 4-band test installation
    • 14-18 Feb: Meerkat visitors (Operations)
    • 16 Feb: Correlator Fire Suppression tests
    • 07-10 Mar: New Worlds, New Horizons (Santa Fe)
    • 07 May - 15 Jun: 4-band science observing
  • Staff Travel:
  • Visitors:
    • 05 Jan - 06 Apr: John Cannon sabbatical at DSOC
    • 22 Feb - 08 Mar: Sonja Vrcic (Correlator)
    • 28 Feb - 28 Mar: Bob Sault

Commissioning priorities for the week (McMullin, Chandler)

Single maintenance day week:
  • Mon
    • 1030-1700: Testing/software development time
    • 1800-0900: Science operations
  • Tue
    • 0900-1700: Testing/software development time
    • 1800-0630: Science operations
  • Wed: Maintenance day
    • 0630-1700: Maintenance operations
    • 1800-0900: Science operations
  • Thu, Fri:
    • 0900-1700: Testing/software development time
    • 1800-0900: Science operations
  • Sat-Sun
    • Science operations + planned testing

Issues for Science Observing:
  • Correlator: fshift sub-band anti-aliasing causing phantom spectral lines when using recirculation for narrow bands (only worry if Michael has to prep your file)
  • Correlator: Missing/mislabeled scans in data sets (small percentage but recurring).
  • Correlator: time mislabeling in SDM data (non-synchronous operation between correlator dumps and CBE can require additional 1/2 integration time padding for online flags);fix available pending testing verification.
  • Construction: RFI sweep at Ku, K, Ka, Q pre-requisite for high pass filters for use with L301.
  • Calibration: Switched power heuristics for reduction.
  • Array/Construction: Dhawan/Rupen: RF switch issues; new set coming to be tested.
  • Correlator: Delay rocking across bands.
  • Correlator: Subset of BlBs misbehaving (only affects hardware stacking modes)
  • Antenna/High Freq: Reference pointing accuracy (C, X, Ku).
  • Array: Amplitude wobbles seen on subset of antennas; phase jumps at scan boundaries also seen on subset of antennas (Ka, Q bands only thus far).

Hardware status and updates (Langley, Perley, Rupen)

  • Array (Perley/Langley):
  • WIDAR (Rupen):
    • Notes: Issues with baseline boards; rebooted and appear to be working but problems from last night are not well understood (suspicion of some versioning issues with CMIB code changes). Plan to resume testing today following meeting, with focus on phased array and low band testing beyond the overall stability investigation.
  • Computing (Robnett):

Software status and updates (Butler)

  • Links:
  • OPT 1.08 pending (22 Feb now); Executor update pending (likely pushed to next week)
  • CASA Update (new functionality available/issues/enhancements):Notes (Miriam): For those using switched power, need asdm2MS (not task-based importxxx) with ignoretime option to get the CalDevice table filled. Missing scan issue still being investigated, tentatively associated with a lustre upgrade (not likely) - talk to Miriam if you see this (to modify the .casarc file); active version uses caching and should be pushed into test soon. Nick Elias has a multi-band delay script available; talk to him/Miriam to test this out; pending this it will be migrated to the public script area.
  • AIPS Update (new functionality available/issues/enhancements): Notes:
  • SDM Update: Notes:SDM FBT news; Bryan/Michael will be reps on this (along with Jeff); requested changes for EVLA involve the Ephemeris table (for planetary) and 3 tables for astrometry (Calc/interferometry tables).

Commissioning Reports

Group updates (User Support, High Frequency, Low Frequency, Spectral Line, Phased Array, Low-band, Algorithm (autoflag))

Commissioning Group updates and Test Planning

  • NOTE: Please cc me on submitted tests so that it can be assigned the appropriate priority!


Individual reports (updates)

  • White: Found issue recurring on a specific channel (Chandler also saw); consistent with internal RFI. Otherwise good.
  • Leroy: Fresh data from the weekend being looked at.
  • Brogan/Hunter: Pending baseline determinations for reduction.


Demo Science Status


  • Shift log tool - a facility for recording issues with the array, allowing both operator/scientist entry of issues and retrieval based on date ranges, antenna/if/etc selection. Not clear how to merge with the existing databases (non-public) for engineering and the correlator. Requirements being reviewed and discussion underway but no quick fix here (though the existing 'spreadsheet' approach is clearly not working adequately). Michael noted that part of Sonja's visit next week will work on alerts within the correlator (better notification on problems in the data will certainly help this problem).
  • Claire noted that there is a meeting at 10am Friday in AOC317 on data distribution for anyone interested.

Misc Actions

  • [Frail/Butler]: Enable restrictions based on proximity to sun; review OST weather trending.
  • [McMullin/Frail]: %REDWrite up disposition of RR items.
  • [McMullin]: Review/recover metrics for Priority 0 items.
  • [Perley/McMullin]: SPW/IF/BB lexicon and mapper
  • Development of WIDAR guidelines document (Michael)
  • List of known problems (can be used by OSRO commissioning team (van Moorsel)
      • Med Priority [TBD]: Added. Phase change when frequency changes (needs to be tested).
      • Med Priority [TBD]: Anti-aliasing does not behave as advertised. - Sault visit * Low Priority [TBD]: Phase (and amp) wobbles; long history.fshift investigation.
-- JosephMcMullin - 2011-02-14
Topic revision: r4 - 2011-02-15, JosephMcMullin
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback